# Dynamical Quantum Phase Transition and Thermal Equilibrium in the Lattice Thirring Model

### v1

@@ V.2 String correlator and DQPT
Some DQPTs have been connected to equilibrium phase diagrams [2, 3, 24, 75]. In Ref. [41], we identified the ground state phase structure by analyzing the behavior of the string correlator (8) (see Sec. III). Namely, the zero-temperature gapless phase is characterized by the string correlator decaying to zero at large distances ($C=0$), while the gapped phase evinces a decay of this correlator to a non-zero value ($C>0$). In the present work, we use similar methodology to connect the occurrence of DQPT to the thermal string correlator, pertaining to the effective temperature of the initial state. As in Ref. [41], we fit the distance-dependent string correlators with the power-exponential ansatz given by Eq. (9). To establish the connection to DQPT, we note that, keeping the mass fixed, $m_{1}a=m_{2}a=0.1$, and varying $\Delta_{1}$, each value of the coupling corresponds to a different effective $\beta$, as illustrated in Fig. 3a (for $\Delta_{2}=-0.9$) and Fig. 3b (for $\Delta_{2}=0.9$).
@@
  • for $\beta\leq 1.563$ ($\Delta_{1}\leq-0.4$), the constant $C$ is consistent with zero (at $\beta=1.563$, $C=0.000004(4)$) and DQPT is observed,
  • for $\beta\geq 1.712$ ($\Delta_{1}\geq-0.3)$, a non-zero constant emerges (at $\beta=1.712$, $C=0.000046(10)$) and we find no DQPT.
The observed correspondence pertains to DQPTs pictured in the lowest branch of Fig. 5b, explicitly marked with circles when $m_{1}a=m_{2}a$. Thus, this kind of DQPT is related to a substantial change in the properties of the thermal state, characterized by a different value of $C$. The former/latter may be the thermal counterpart of the zero-temperature critical/gapped phase, but a detailed characterization of finite temperature phases of the model is beyond the scope of our work.
Figure 6: Value of the constant $C$, defined in Eq. (9), that characterizes the long distance behavior of the string correlator (8), for the thermal equilibrium state of the Hamiltonian $(m_{2}a,\Delta_{2})=(0.1,0.9)$ as a function of the inverse temperature $\beta$. The grey dashed line in the inset denotes the critical inverse temperature $\beta_{c}$.
The picture is somewhat different for $\Delta_{2}=-0.9$. In this case, $C=0$ at all temperatures, regardless of the occurrence of DQPT, found at $\beta\lesssim 0.34$ (see Fig. 5a). Note, however, that none of the DQPTs represented in Fig. 5a occurrs at $m_{1}a=m_{2}a$, i.e. a prerequisite for DQPT at $\Delta_{2}=-0.9$ is to have a quench in fermion mass. Thus, the analyzed thermal string correlator is not related in any way to the presence of a DQPT for the final Hamiltonian $(m_{2}a,\,\Delta_{2})$=(0.1,-0.9). However, it cannot be excluded that this DQPT is related to a phase transition between other types of phases, characterized by a different observable. We leave a more detailed investigation of finite-temperature phases and their potential relation to DQPTs for future work.
Ultimately, we conclude that there is no direct correspondence between the presence of a DQPT and the phase structure probed by the decay of the string correlator. It is plausible that the DQPTs observed for $m_{2}a=0.1$ and $\Delta_{2}=-0.9$ or $\Delta_{2}=0.9$ are of a different type, with only the latter related to the large-distance properties described by the string correlator.
@@ VI Conclusion
We have identified a threshold in the energy density of the initial state below which DQPTs do not occur, at least during the finite times reachable by our simulations. The threshold is observed both for quenches in the gapless and the gapped phases. The energy density of each initial state corresponds to an effective temperature: that of the thermal equilibrium state with the same energy per spin, restricted to the sector of zero magnetization. This thermal state is the one expected to describe the equilibrium values of local observables after the quench in the long time limit, since energy and magnetization are the only local conserved quantities in the problem.
We have analyzed the position of each observed DQPT (or non-analyticity of the Loschmidt rate) in the time-effective-temperature plane. This reveals structures of non-analyticities reminiscent of the zeros in the complex-time plane of the Loschmidt rate defined for a fixed initial state. Our results show that DQPTs for different initial states lie in multiple well-defined branches. Initial states in different phases can appear on the same branch, which thus is not related to the zero-temperature phase diagram. All DQPT branches start at some minimal effective temperature. In most cases, the lowest temperature edge of a branch corresponds to a smooth merge with another one. However, in the case of the quench with the gapped Hamiltonian, we identify a particular branch, with DQPTs happening at the shortest times, with a sharp end at $\beta_{c}\sim 1.7$. We have found that this energy threshold corresponds very accurately with a substantial change in the properties of the thermal equilibrium states at the corresponding effective temperature. More concretely, below this temperature, a certain string correlator attains a non-zero value at long distances, a property that also characterizes the ground state in the gapped phase. For this branch, we have thus established a relation between the presence of DQPTs and the equilibrium phase diagram at finite T. We have not found a similar correspondence for the smooth thresholds of branches that merge into each other.
Our study thus suggests that observing the dynamics and analyzing the DQPTs from different initial states can probe properties of the finite temperature phase diagram, in particular when a sharp threshold is found for the occurrence of these non-analyticities. Since our simulations are limited to finite time windows, due to the accumulation of truncation error, and we have not mapped out the complete thermal equilibrium properties at all temperatures, further investigation is needed to determine whether the other observed branches correspond also to some quantifiable equilibrium property. It is also interesting to investigate whether similar effects appear in other spin models, and the precise correspondence between Fisher zeros and the DQPT lines in our time-effective temperature representation.
@@ Acknowledgements.
The authors thank I. McCulloch, L. Tagliacozzo and M. Fagotti for very useful discussions. This work was partly supported by the DFG (German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868, and Research Unit FOR 5522 (grant nr. 499180199); and by the EU-QUANTERA project TNiSQ (BA 6059/1-1), as well as Taiwanese NSTC grants, 110-2112-M-002-034-MY3, 112-2112-M-A49-021-MY3, 112-2119-M-007-008 and 113-2119-M-007-013. Numerical computations were performed on HPC facilities at National Taiwan University and National Yang Ming Chiao Tung University. The data and code to generate the figures are available at https://github.com/hunghaoti/Thirring_data.
@@ Appendix A Penalty term and Matrix Product Operator
Since the $z-$component of the total spin corresponds to the total fermion number in the Thirring model, and we are only interested in the zero-charge sector. To enforce this, we introduce a penalty term in the Hamiltonian,
@@
Here, $\lambda$ should be sufficiently large (100 in this work) to ensure that the ground state obtained via a variational search is in the sector of vanishing total $S^{z}$ [76], that is $\langle{S_{tot}^{z}}\rangle=\langle{\sum_{n}{S_{n}^{z}}}\rangle=0$. This enables us to interpret our results in terms of the dual SG theory and the XY model. However, the penalty term will not be turned on during the real time evolution simulations since $\langle S_{tot}^{z}\rangle$ is a conserved quantity. In our simulations, this quantity is well-preserved, i.e., $\langle S_{tot}^{z}(t)\rangle=0$.
In TN algorithms, it is useful to express the Hamiltonian as the matrix product operator(MPO). The Hamiltonian of Thirring model in Eq. (18) can be expressed as an MPO:
@@
where $\beta_{n}=\Delta+(-1)^{n}\tilde{m_{0}}a,\,\alpha=\frac{1}{4}\left(\lambda+\Delta\right)$. The $\beta_{n}$ contains the staggered term. To simplify the simulation, we also combine two-site MPOs into one-site, as we did with the uMPS, represented as $W_{1}W_{2}\rightarrow W$.
@@ Appendix B Brief introduction of VUMPS and TDVP
For real-time evolution, we choose the TDVP algorithm, which preserves conserved quantities well [72]. TDVP corresponds to a projection of the evolution state $\ddot{\text{o}}$dinger equation onto the tangent space of the uMPS manifold $\mathcal{M}$ at the current state (see Fig. 8, ). We can solve the time derivative of the uMPS $|\Psi(A)\rangle$ through the TDVP equation:

### v2

@@ V.2 String correlator and DQPT
Some DQPTs have been connected to equilibrium phase diagrams [2, 3, 24, 77]. In Ref. [41], we identified the ground-state phase structure by analyzing the behavior of the string correlator (III) (see Sec. III). Namely, the zero-temperature gapless phase is characterized by the string correlator decaying to zero at large distances ($C=0$), while the gapped phase evinces a decay of this correlator to a nonzero value ($C>0$). In the present work, we use similar methodology to connect the occurrence of DQPT to the thermal string correlator, pertaining to the effective temperature of the initial state. As in Ref. [41], we fit the distance-dependent string correlators with the power-exponential ansatz given by Eq. (9). To establish the connection to DQPT, we note that, keeping the mass fixed, $m_{1}a=m_{2}a=0.1$, and varying $\Delta_{1}$, each value of the coupling corresponds to a different effective $\beta$, as illustrated in Fig. 4(a) (for $\Delta_{2}=-0.9$) and Fig. 4(b) (for $\Delta_{2}=0.9$).
@@
     (i)  for $\beta\leq 1.563$ ($\Delta_{1}\leq-0.4$), the constant $C$ is consistent with zero [at $\beta=1.563$, $C=0.000004(4)$] and DQPT is observed,
     (ii)  for $\beta\geq 1.712$ ($\Delta_{1}\geq-0.3)$, a nonzero constant emerges [at $\beta=1.712$, $C=0.000046(10)$] and we find no DQPT.
The observed correspondence pertains to DQPTs pictured in the lowest branch of Fig. 6(b), explicitly marked with circles when $m_{1}a=m_{2}a$. Thus, this kind of DQPT is related to a substantial change in the properties of the thermal state, characterized by a different value of $C$. The former/latter may be the thermal counterpart of the zero-temperature critical/gapped phase, but a detailed characterization of finite-temperature phases of the model is beyond the scope of our work.
Figure 7: Value of the constant $C$, defined in Eq. (9), that characterizes the long-distance behavior of the string correlator (III) for the thermal equilibrium state of the Hamiltonian $(m_{2}a,\Delta_{2})=(0.1,0.9)$ as a function of the inverse temperature $\beta$. The gray dashed line in the inset denotes the critical inverse temperature $\beta_{c}$.
The picture is somewhat different for $\Delta_{2}=-0.9$. In this case, $C=0$ at all temperatures, regardless of the occurrence of DQPT, found at $\beta\lesssim 0.34$ [see Fig. 6(a)]. Note, however, that none of the DQPTs represented in Fig. 6(a) occur at $m_{1}a=m_{2}a$, i.e., a prerequisite for DQPT at $\Delta_{2}=-0.9$ is to have a quench in fermion mass. Thus, for the final Hamiltonian in the critical phase $(m_{2}a,\,\Delta_{2})=(0.1,\,-0.9)$, the analyzed string correlator does not exhibit a similar correlation to the presence of DQPTs as the one observed for the quench into the gapped Hamiltonian. This might be understood because the asymptotic value of the string correlator detects the order present in the ground state of the gapped phase, which is destroyed as the temperature (energy) increases, whereas the ground state in the critical case does not exhibit such order. On the other hand, the DQPT threshold detected by the correlator in the gapped quench corresponds to the abrupt edge of the lower DQPT branch in Fig. 6(b), a feature that does not appear either in the corresponding Fig. 6(a) for the critical quench. Although we conclude that the thermal string correlators defined in Eq. (III) are not universal indicators for the possibility of DQPTs, a more detailed investigation is needed to fully characterize which DQPTs are associated to changes in these quantities and to determine whether other thermal observables can detect the different boundaries of DQPT branches observed in Fig. 6. We thus leave the investigation of finite-temperature phases and their potential relation to DQPTs for future work.
@@ VI Conclusion
We have identified a threshold in the energy density of the initial state below which DQPTs do not occur, at least during the finite times reachable by our simulations. The threshold is observed both for quenches in the gapless and the gapped phases. The energy density of each initial state corresponds to an effective temperature: that of the thermal equilibrium state with the same energy per spin, restricted to the sector of zero magnetization. This thermal state is the one expected to describe the equilibrium values of local observables after the quench in the long-time limit, since energy and magnetization are the only local conserved quantities in the problem.
We have analyzed the position of each observed DQPT (or nonanalyticity of the Loschmidt rate) in the time-effective-temperature plane. This reveals structures of nonanalyticities reminiscent of the zeros in the complex-time plane of the Loschmidt rate defined for a fixed initial state. Our results show that DQPTs for different initial states lie in multiple well-defined branches. Initial states in different phases can appear on the same branch, which thus is not related to the zero-temperature phase diagram. All DQPT branches start at some minimal effective temperature. In most cases, the lowest temperature edge of a branch corresponds to a smooth merge with another one. However, in the case of the quench with the gapped Hamiltonian, we identify a particular branch, with DQPTs happening at the shortest times, with a sharp end at $\beta_{c}\sim 1.7$. We have found that this energy threshold corresponds very accurately with a substantial change in the properties of the thermal equilibrium states at the corresponding effective temperature. More concretely, below this temperature, a certain string correlator attains a nonzero value at long distances, a property that also characterizes the ground state in the gapped phase. For this branch, we have thus established a relation between the presence of DQPTs and the equilibrium phase diagram at finite $T$. We have not found a similar correspondence for the smooth thresholds of branches that merge into each other.
Our study thus suggests that observing the dynamics and analyzing the DQPTs from different initial states can probe properties of the finite-temperature phase diagram, in particular when a sharp threshold is found for the occurrence of these nonanalyticities. Since our simulations are limited to finite-time windows, due to the accumulation of truncation error, and we have not mapped out the complete thermal equilibrium properties at all temperatures, further investigation is needed to determine whether the other observed branches correspond also to some quantifiable equilibrium property. It is also interesting to investigate whether similar effects appear in other spin models, and the precise correspondence between Fisher zeros and the DQPT lines in our time-effective temperature representation.
@@ Acknowledgements.
The authors thank I. McCulloch, L. Tagliacozzo, M. Fagotti, L. Barbiero, and N. Cuzzuol for very useful discussions. This work was partly supported by the DFG (German Research Foundation) under Germany’s Excellence Strategy – EXC-2111 – 390814868 and Research Unit FOR 5522 (Grant No. 499180199) and by the EU-QUANTERA project TNiSQ (BA 6059/1-1), as well as Taiwanese NSTC Grants No. 110-2112-M-002-034-MY3, No. 112-2112-M-A49-021-MY3, No. 112-2119-M-007-008, and No. 113-2119-M-007-013. Numerical computations were performed on HPC facilities at National Taiwan University and National Yang Ming Chiao Tung University.
Data availability
The data and code to generate the figures are available in Github [78].
@@ Appendix A Penalty Term and Matrix Product Operator
Since the $z$ component of the total spin corresponds to the total fermion number in the Thirring model, and we are only interested in the zero-charge sector. To enforce this, we introduce a penalty term in the Hamiltonian,
@@
Here $\lambda$ should be sufficiently large (100 in this work) to ensure that the ground state obtained via a variational search is in the sector of vanishing total $S^{z}$ [79], that is, $\langle{S_{tot}^{z}}\rangle=\langle{\sum_{n}{S_{n}^{z}}}\rangle=0$. This enables us to interpret our results in terms of the dual SG theory and the XY model. However, the penalty term will not be turned on during the real-time evolution simulations since $\langle S_{tot}^{z}\rangle$ is a conserved quantity. In our simulations, this quantity is well preserved, i.e., $\langle S_{tot}^{z}(t)\rangle=0$.
In TN algorithms, it is useful to express the Hamiltonian as the MPO. The Hamiltonian of Thirring model in Eq. (18) can be expressed as an MPO:
@@
where $\beta_{n}=\Delta+(-1)^{n}\tilde{m_{0}}a,\,\alpha=\frac{1}{4}\left(\lambda+\Delta\right)$. The $\beta_{n}$ contains the staggered term. To simplify the simulation, we also combine two-site MPOs into one site, as we did with the uMPS, represented as $W_{1}W_{2}\rightarrow W$.
@@ Appendix B Brief Introduction of VUMPS and TDVP
For real-time evolution, we choose the TDVP algorithm, which preserves conserved quantities well [73]. TDVP corresponds to a projection of the evolution state $\ddot{\text{o}}$dinger equation onto the tangent space of the uMPS manifold $\mathcal{M}$ at the current state (see Fig. 9). We can solve the time derivative of the uMPS $|\Psi(A)\rangle$ through the TDVP equation: